# GUIDE: Guidance-based Incremental Learning with Diffusion Models

### 06 Mar 2024 v1

@@ Abstract
We introduce GUIDE, a novel continual learning approach that directs diffusion models to rehearse samples at risk of being forgotten. Existing generative strategies combat catastrophic forgetting by randomly sampling rehearsal examples from a generative model. Such an approach contradicts buffer-based approaches where sampling strategy plays an important role. We propose to bridge this gap by integrating diffusion models with classifier guidance techniques to produce rehearsal examples specifically targeting information forgotten by a continuously trained model. This approach enables the generation of samples from preceding task distributions, which are more likely to be misclassified in the context of recently encountered classes. Our experimental results show that GUIDE significantly reduces catastrophic forgetting, outperforming conventional random sampling approaches and surpassing recent state-of-the-art methods in continual learning with generative replay.
Machine Learning, ICML
Figure 1: Rehearsal sampling in GUIDE. We guide the denoising process of a diffusion model trained on the previous task (blue) toward classes from the current task (orange). The replay samples, highlighted with blue borders, share features with the examples from the current task, which may be related to color (e.g., blue cars when guided to airplanes) or other characteristics (e.g., cars on a grassy surface when guided to horses). Continual training of a classifier on such samples positioned near its decision boundary successfully mitigates catastrophic forgetting.
@@ 1 Introduction
A typical machine learning pipeline involves training a model on a static dataset and deploying it in a task with a similar data distribution. This assumption frequently proves impractical in real-world scenarios, where models encounter a constantly evolving set of objectives. To address this issue, Continual Learning (CL) methods try to accumulate knowledge from separate tasks while overcoming limited knowledge transfer and catastrophic forgetting (French, 1999).
An effective CL strategy is to rehearse examples from previous tasks by storing them in a memory buffer and mixing them with samples from the current task. This, however, requires a well-designed sampling strategy, therefore many successful variants have been proposed, including: approximating the true data distribution with stored samples (Rebuffi et al., 2017; Isele & Cosgun, 2018), augmenting past samples with features observed in the current task (Kumari et al., 2022), or hard negative mining (Jin et al., 2021). Despite offering decent performance, these methods are unsuitable for certain applications, e.g., due to limited scalability, memory constraints, or privacy restrictions. To mitigate this issue, Shin et al. (2017) introduced Deep Generative Replay (DGR) that substitutes the memory buffer with a Generative Adversarial Networks (Goodfellow et al., 2014), which was later extended with various generative models (Van de Ven et al., 2020; Scardapane et al., 2020; Gao & Liu, 2023). Although generative rehearsal approaches mitigate the limitations of buffer-based methods, relatively little attention was paid to understanding and controlling what samples they generate, simply relying on random generations. This contradicts the importance of sampling strategies proposed for buffer-based techniques. In this work, we propose to bridge this gap and introduce GUIDE: a new continual learning method with Denoising Deep Probabilistic Models (DDPM)  (Sohl-Dickstein et al., 2015; Ho et al., 2020). We first show that we can use the classifier guidance technique (Dhariwal & Nichol, 2021) to steer the backward diffusion process towards classes not present in the diffusion’s training data. Then, motivated by the observation that examples located close to the decision boundary have a higher chance of being forgotten (Toneva et al., 2018), we propose to use the previous observation in generative rehearsal and guide the diffusion model trained on data from preceding tasks towards recently encountered classes. The visualization of this idea is presented in Fig. 1.
@@ 2.1 Continual Learning
Continual Learning (CL) methods aim to mitigate catastrophic forgetting – a situation where deep neural networks trained on a sequence of tasks completely and abruptly forget previously learned information upon retraining on a new task. Recent methods can be organized into three main families. Regularization methods (Kirkpatrick et al., 2017; Zenke et al., 2017; Li & Hoiem, 2017) identify the most important parameters and try to slow down their changes through regularization. Architectural approaches (Rusu et al., 2016; Yoon et al., 2018; Mallya & Lazebnik, 2018; Mallya et al., 2018) change the structure of the model for each task. Rehearsal methods replay data samples from previous tasks and train the model on a combination of data samples from previous and current tasks. In the simplest setup, a memory buffer is used to store previous samples for simple experience replay (Prabhu et al. (2020), Lopez-Paz & Ranzato (2017), Aljundi et al. (2019)) or store exemplars to prevent forgetting in class incremental learning (Rebuffi et al., 2017; Chaudhry et al., 2018; Wu et al., 2019; Hou et al., 2019; Belouadah & Popescu, 2019; Castro et al., 2018). Some methods instead of using directly exemplars, stores data representations of a previous task in different forms, e.g., mnemonics - optimized artificial samples (Liu et al., 2020b), distilled datasets (Wang et al., 2018; Zhao & Bilen, 2021; Zhao et al., 2021), or addressable memory structure (Deng & Russakovsky, 2022).
@@
Because of the limitations of buffer-based approaches related to the constantly growing memory requirements and privacy issues, Shin et al. (2017) introduced Deep Generative Replay, where a Generative Adversarial Network (GAN) is used to learn data distribution from the previous tasks in order to generate rehearsal samples for the continually trained classifier. A similar approach is further extended to different model architectures like Variational Autoencoders (VAE) (van de Ven & Tolias, 2018; Nguyen et al., 2018), normalizing flows (Scardapane et al., 2020) or Gaussian Mixture Models (Rostami et al., 2019).
On top of those baseline approaches, Ramapuram et al. (2020) introduce a method that benefits from the knowledge distillation technique in VAE training in CL setup, while Wu et al. (2018) introduce Memory Replay GANs (MeRGANs) and describe two approaches to prevent forgetting - by joint retraining and by aligning the replay samples.
Instead of replaying images, several approaches propose to rehearse internal data representations instead, e.g., Brain-Inspired Replay (BIR) (Van de Ven et al., 2020) with an extension to Generative Feature Replay (GFR) (Liu et al., 2020a), where rehearsal is combined with features distillation. Kemker & Kanan (2018) divide feature rehearsal into short and long-term parts.
@@ 2.2 Guided image generation in diffusion models
Besides conditioning, controlling diffusion model outputs can be achieved by modifying the process of sampling that incorporates additional signals from the guidance function. Classifier guidance introduced by Dhariwal & Nichol (2021) employs a trained classifier model to steer the generation process via gradients of a specified loss function, typically assessing the alignment of generated images with specific classes. This concept is extended by Bansal et al. (2023) to include any off-the-shelf model in guidance and further applied by Augustin et al. (2022) for explaining the decisions of the classifier with generated counterfactual examples. Contrasting with these methods, Epstein et al. (2023) introduce self-guidance based on the internal representations of the diffusion model, while Ho & Salimans (2022) introduce classifier-free guidance, achieving results akin to classifier-based approaches by joint training of unconditional and conditional diffusion models.
Continual learning with diffusion models
Diffusion models excel in generative tasks, surpassing VAEs (Kingma & Welling, 2014) and GANs, yet their adoption in CL remains limited. Masip et al. (2023) introduce an approach that distills the reverse diffusion process and applies this idea to Generative Replay (GR). Smith et al. (2023) explore continual training of text-to-image diffusion models, while Zając et al. (2023) reviews diffusion models in CL, discussing the dynamics of their forgetting. Deep Diffusion-based Generative Replay (DDGR) (Gao & Liu, 2023) uses a diffusion model in a generative rehearsal method and benefits from a classifier pretrained on previous tasks to synthesize high-quality replay samples. Class-Prototype Conditional Diffusion Model (CDPM) (Doan et al., 2023) further enhances the replay samples quality by conditioning the diffusion model on learnable class-prototypes. In this work, we further extend those ideas and show that classifier guidance might be used not only to generate high-quality samples but also to introduce the desired characteristics of rehearsal samples.
@@ 3.1 Diffusion models
In this work, we follow the formulation of DDPM presented by Dhariwal & Nichol (2021).
Let $q(\mathbf{x}_{0})$ be the real data distribution. The forward diffusion process $q$ gradually adds Gaussian noise to each data sample $\mathbf{x}_{0}\sim q(\mathbf{x}_{0})$ over a sequence of $T$ steps according to the variance schedule $\{\beta_{t}\in(0,1)\}_{t=1}^{T}$:
$q(\mathbf{x}_{t}|\mathbf{x}_{t-1})=\mathcal{N}(\mathbf{x}_{t};\sqrt{1-\beta_{t}}\mathbf{x}_{t-1},\beta_{t}\mathbf{I}).$ (1)
This process produces a sequence of noisy samples $\mathbf{x}_{1},\dots,\mathbf{x}_{T}$, where $\mathbf{x}_{T}$ is pure Gaussian noise.
The reverse diffusion process $p$ can be modeled as a diagonal Gaussian:
@@
Sampling from DDPM is expensive as it requires denoising the sample sequentially from $\mathbf{x}_{T}\sim\mathcal{N}(0,\mathbf{I})$ through many steps $T$ up until $\mathbf{x}_{0}$. Thus, in this work, we benefit from different definitions of a reverse diffusion process, known as Denoising Diffusion Implicit Models (DDIM) introduced by Song et al. (2020), where at each time step $t$ we first predict the denoised sample at $t=0$ as:
$\hat{\mathbf{z}}_{0}(\mathbf{x}_{t})=\frac{\mathbf{x}_{t}-\sqrt{1-\alpha_{t}}\epsilon_{\theta}(\mathbf{x}_{t},t)}{\sqrt{\alpha_{t}}}.$ (4)
Then we can generate $\mathbf{x}_{t-1}$ from $\mathbf{x_{t}}$ using following formula:
$\mathbf{x}_{t-1}=\sqrt{\alpha_{t-1}}\hat{\mathbf{z}}_{0}(\mathbf{x}_{t})+\sqrt{1-\alpha_{t-1}}\epsilon_{\theta}(\mathbf{x}_{t},t).$ (5)
DDIM allows us to sample the diffusion model using fewer steps in the reverse process, reducing the inference time significantly.
@@ 3.2 Classifier guidance
Classifier guidance, introduced by Dhariwal & Nichol (2021), enables steering the backward diffusion process by combining the intermediate diffusion steps of conditional or unconditional diffusion with the gradient from the externally trained classifier. To ensure the high quality of generated samples, the original method was introduced with a classifier trained on noised images. However, such an approach can be impractical and lead to limited performance when evaluated on clear data. Building upon work by Bansal et al. (2023), in our method, we adapt the guidance process to utilize a classifier trained only on clean images.
To that end, in guidance process we first predict $\hat{\mathbf{z}}_{0}(\mathbf{x}_{t})$ using diffusion model, as in Eq. 4. We can then modify the prediction of diffusion model $\epsilon_{\theta}(\mathbf{x}_{t},t)$ at each time step $t$ according to:
@@ 3.3 Continual learning with generative replay
In this work, we focus on class-incremental continual learning (van de Ven et al., 2022) of a classifier $f_{\phi}(y|\mathbf{x})$ with generative replay, where we use diffusion model $\epsilon_{\theta}(\mathbf{x}_{t},t,y)$ as a generator of synthetic samples from previous tasks. The data stream comprises of $T$ distinct tasks. Each task $i$ has its associated dataset $\mathcal{D}_{i}$ with collection of $N$ pairs $\{(\mathbf{x}_{i}^{j},y_{i}^{j})\}_{j=1}^{N}$, where $\mathbf{x}_{i}^{j}$ and $y_{i}^{j}$ represent the $j$-th input sample and its corresponding label, respectively.
@@ 4 Method
This section introduces GUIDE - a novel method designed to mitigate catastrophic forgetting in a classifier trained in a generative replay scenario with a diffusion model. We benefit from the classifier trained on the current task, referred to as the current classifier, to guide the diffusion model during the generation of rehearsal examples from preceding tasks. With our method, we can generate rehearsal examples close to the classifier’s decision boundary, making them highly valuable to counteract the classifier’s forgetting in class-incremental learning (Toneva et al., 2018; Kumari et al., 2022).
@@ 4.1 Intuition and rationale behind GUIDE
Before moving to the continual-learning setup, we demonstrate the effect of guiding the diffusion model towards classes not included in its training dataset in this section. To that end, we propose a simplified scenario in which we train an unconditional diffusion model exclusively on the first two classes from the CIFAR-10 dataset (airplane and automobile). Additionally, we train two separate classifiers: $f_{\phi_{1}}(y_{1}|\mathbf{x})$, aligned with the diffusion model’s training classes, and $f_{\phi_{2}}(y_{2}|\mathbf{x})$, trained with all data from the CIFAR-10 dataset. Given the model’s unconditional nature, we employ the classifier $f_{\phi_{1}}(y_{1}|\mathbf{x})$ to steer the diffusion denoising process towards one of the classes that it is trained on. Simultaneously, $f_{\phi_{2}}(y_{2}|\mathbf{x})$ guides the generation process to one of the classes from the CIFAR-10 dataset. Formally, we modify the unconditional diffusion sampling process as follows:
$\begin{split}{\hat{\epsilon}}_{\theta}(\mathbf{x}_{t},t)&=\epsilon_{\theta}(\mathbf{x}_{t},t)\\ &+s_{1}\nabla_{\mathbf{x}_{t}}{\ell}\left(f_{\phi_{1}}(y_{1}|\hat{\mathbf{z}}_{0}(\mathbf{x}_{t})),c_{1}\right)\\ &+s_{2}\nabla_{\mathbf{x}_{t}}{\ell}\left(f_{\phi_{2}}(y_{2}|\hat{\mathbf{z}}_{0}(\mathbf{x}_{t})),c_{2}\right),\end{split}$ (9)
where $c_{1}\in y_{1}$ and $c_{2}\in y_{2}$ are CIFAR-10 classes to which we guide with the classifiers $f_{\phi_{1}}$ and $f_{\phi_{2}}$ respectively, with $s_{1}$ and $s_{2}$ as their corresponding classifier scales. In Fig. 2, we present generated samples, with each row showcasing generations towards a different class $c_{2}$.
Figure 2: Samples from the unconditional diffusion model trained only on airplanes and automobiles from the CIFAR-10 dataset. We guide the denoising diffusion process with two separate classifiers: the first one guides towards airplanes (left) or automobiles (right), while the second one to the class presented on the left side of the figure. We obtain samples from the desired class with observable features of classes unknown to the diffusion model, such as the color or the shape.
By integrating guidance from both classifiers, we generate samples from the diffusion model’s training data distribution – all of the examples are either planes or cars, but with visible features from classes $y_{2}$, which are unknown to the diffusion model (e.g., grassy surfaces when guided towards horses, frog-like objects colors, or bird-like shapes and colors).
In this work, we propose to use this observation in the continual training of a classifier, with the distinction that in GUIDE, we only utilize guidance from one currently trained classifier because we sample from a class-conditional diffusion model, eliminating the need for guidance from two classifiers.

### 31 May 2024 v2

@@ Abstract
We introduce GUIDE, a novel continual learning approach that directs diffusion models to rehearse samples at risk of being forgotten. Existing generative strategies combat catastrophic forgetting by randomly sampling rehearsal examples from a generative model. Such an approach contradicts buffer-based approaches where sampling strategy plays an important role. We propose to bridge this gap by incorporating classifier guidance into the diffusion process to produce rehearsal examples specifically targeting information forgotten by a continuously trained model. This approach enables the generation of samples from preceding task distributions, which are more likely to be misclassified in the context of recently encountered classes. Our experimental results show that GUIDE significantly reduces catastrophic forgetting, outperforming conventional random sampling approaches and surpassing recent state-of-the-art methods in continual learning with generative replay.
@@ 1 Introduction
A typical machine learning pipeline involves training a model on a static dataset and deploying it in a task with a similar data distribution. This assumption frequently proves impractical in real-world scenarios, where models encounter a constantly evolving set of objectives. To address this issue, Continual Learning (CL) methods try to accumulate knowledge from separate tasks while overcoming limited knowledge transfer and catastrophic forgetting [11]. Replay-based CL strategies employ a memory buffer to retain examples from preceding tasks, which are then mixed with samples from the current task. In order for this approach to be effective, it requires a well-designed sampling strategy such as approximating the true data distribution with stored samples [36, 20], maximizing the diversity of samples in a buffer [53], or hard negative mining [21].
Figure 1: Rehearsal sampling in GUIDE. We guide the denoising process of a diffusion model trained on the previous task (blue) toward classes from the current task (orange). The replay samples, highlighted with blue borders, share features with the examples from the current task, which may be related to characteristics such as color or background (e.g., fishes on a snowy background when guided to snowmobile). Generative rehearsal on such samples positioned near the classifier’s decision boundary successfully mitigates catastrophic forgetting.
Despite offering decent performance, these methods are unsuitable for certain applications, e.g., due to limited scalability, memory constraints, or privacy restrictions. To mitigate this issue, Shin et al. [41] introduced Deep Generative Replay (DGR) that substitutes the memory buffer with a Generative Adversarial Networks (GAN) [13], which was later extended with various generative models [46, 39, 12]. Although generative rehearsal approaches mitigate the limitations of buffer-based methods, relatively little attention was paid to understanding and controlling what samples they generate, simply relying on random generations. This contradicts the importance of sampling strategies proposed for buffer-based techniques.
In this work, we propose to bridge this gap and introduce GUIDE: a new continual learning method with Denoising Deep Probabilistic Models (DDPM)  [42, 18]. We first demonstrate that we can use the classifier guidance technique [8] to steer the backward diffusion process towards classes not present in the training set of a diffusion model. Then, additionally motivated by the finding that examples located close to the decision boundary are more likely to be forgotten [44], we guide the diffusion model trained on data from preceding tasks towards recently encountered classes in generative rehearsal scenario. The visualization of this idea is presented in Fig. 1.
@@ 2.1 Continual Learning
Continual learning methods aim to mitigate catastrophic forgetting – a phenomenon where deep neural networks trained on a sequence of tasks completely and abruptly forget previously learned information upon retraining on a new task. Recent methods can be organized into three main families. Regularization methods [24, 54, 28] identify the most important parameters and try to slow down their changes through regularization. Architectural approaches [38, 52, 31, 32] change the structure of the model for each task. Rehearsal methods replay data samples from previous tasks and train the model on a combination of data samples from previous and current tasks. In the most straightforward rehearsal approach, a memory buffer is used to store exemplars from previous tasks [34, 36, 5, 51, 19, 3, 4]. Some methods, instead of directly using exemplars, stores data representations from previous tasks in different forms, e.g., mnemonics - optimized artificial samples [30], distilled datasets [48, 55, 56], or addressable memory structure [7].
@@
Because of the limitations of buffer-based approaches related to the constantly growing memory requirements and privacy issues, Shin et al. [41] introduce Deep Generative Replay, where a GAN is used to generate rehearsal samples from previous tasks for the continual training of a classifier. A similar approach is further extended to different model architectures like Variational Autoencoders (VAE) [45, 33], normalizing flows [39] or Gaussian Mixture Models [37].
On top of those baseline approaches, Ramapuram et al. [35] introduce a method that benefits from the knowledge distillation technique in VAE training in CL setup, while Wu et al. [50] introduce Memory Replay GANs (MeRGANs) and describe two approaches to prevent forgetting - by joint retraining and by aligning the replay samples.
Instead of replaying images, several approaches propose to rehearse internal data representations instead, e.g., Brain-Inspired Replay (BIR) [46] with an extension to Generative Feature Replay (GFR) [29], where the rehearsal is combined with features distillation. Kemker and Kanan [22] divide feature rehearsal into short and long-term parts.
@@ 2.2 Guided image generation in diffusion models
Besides conditioning, controlling diffusion model outputs can be achieved by modifying the process of sampling that incorporates additional signals from the guidance function. Classifier guidance introduced by Dhariwal and Nichol [8] employ a trained classifier model to steer the generation process via gradients of a specified loss function, typically assessing the alignment of generated images with specific classes. This concept is extended by Bansal et al. [2] to include any off-the-shelf model in guidance and further applied by Augustin et al. [1] for explaining the decisions of the classifier with generated counterfactual examples. Contrasting with these methods, Epstein et al. [10] introduce self-guidance based on the internal representations of the diffusion model, while Ho and Salimans [17] introduce classifier-free guidance, achieving results akin to classifier-based approaches by joint training of unconditional and conditional diffusion models.
2.3 Continual learning with diffusion models
Diffusion models excel in generative tasks, surpassing VAEs [23] and GANs, yet their adoption in CL remains limited. Deep Diffusion-based Generative Replay (DDGR) [12] uses a diffusion model in a generative rehearsal method and benefits from a classifier pretrained on previous tasks to synthesize high-quality replay samples. Class-Prototype Conditional Diffusion Model (CDPM) [9] further enhances the replay samples quality by conditioning the diffusion model on learnable class-prototypes. In this work, we further extend those ideas and show that classifier guidance might be used not only to generate high-quality samples but also to introduce the desired characteristics of rehearsal samples.
@@ 3.1 Diffusion models
Let $q(\mathbf{x}_{0})$ be the real data distribution. The forward diffusion process $q$ gradually adds Gaussian noise to each data sample $\mathbf{x}_{0}\sim q(\mathbf{x}_{0})$ over a sequence of $T$ steps according to the variance schedule $\{\beta_{t}\in(0,1)\}_{t=1}^{T}$. If we additionally define $\alpha_{t}=1-\beta_{t}$ and $\bar{\alpha}_{t}=\prod_{i=1}^{t}\alpha_{i}$, then:
$q(\mathbf{x}_{t}|\mathbf{x}_{t-1})=\mathcal{N}(\mathbf{x}_{t};\sqrt{\alpha_{t}}\mathbf{x}_{t-1},(1-\alpha_{t})\mathbf{I}).$ (1)
This process produces a sequence of noisy samples $\mathbf{x}_{1},\dots,\mathbf{x}_{T}$, where $\mathbf{x}_{T}$ is pure Gaussian noise. The reverse diffusion process $p$ can be modeled as a diagonal Gaussian:
@@
Sampling from DDPM requires denoising the sample sequentially from $\mathbf{x}_{T}\sim\mathcal{N}(0,\mathbf{I})$ $T$ steps up until $\mathbf{x}_{0}$. In this work, we benefit from the speed-up technique known as Denoising Diffusion Implicit Models (DDIM) introduced by Song et al. [43], where at each time step $t$, we first predict the denoised sample at $t=0$:
$\hat{\mathbf{z}}_{0}(\mathbf{x}_{t})=\frac{\mathbf{x}_{t}-\sqrt{1-\bar{\alpha}_{t}}\epsilon_{\theta}(\mathbf{x}_{t},t)}{\sqrt{\bar{\alpha}_{t}}},$ (4)
then we calculate $\mathbf{x}_{t-1}$ based on this prediction. This allows us to sample using fewer steps in the reverse process, reducing the inference time significantly.
@@ 3.2 Classifier guidance
Classifier guidance [8] enables steering the backward diffusion process by combining the intermediate denoising steps of conditional or unconditional diffusion with the gradient from the externally trained classifier. To ensure the high quality of generated samples, the original method was introduced with a classifier trained on noised images. However, such an approach can be impractical and lead to limited performance.
Building upon the work by Bansal et al. [2], in our method, we adapt the guidance process to utilize a classifier trained only on clean images. To that end, in the guidance process, we first predict denoised image $\hat{\mathbf{z}}_{0}(\mathbf{x}_{t})$, as in Eq. 4. We can then modify the prediction of diffusion model $\epsilon_{\theta}(\mathbf{x}_{t},t)$ at each time step $t$ according to:
@@ 3.3 Continual learning with generative replay
In this work, we focus on class-incremental continual learning [47] of a classifier $f_{\phi}(y|\mathbf{x})$ with generative replay, where we use diffusion model $\epsilon_{\theta}(\mathbf{x}_{t},t,y)$ as a generator of synthetic samples from previous tasks. The data stream comprises of $T$ distinct tasks. Each task $i$ has its associated dataset $\mathcal{D}_{i}$ with collection of $N$ pairs $\{(\mathbf{x}_{i}^{j},y_{i}^{j})\}_{j=1}^{N}$, where $\mathbf{x}_{i}^{j}$ and $y_{i}^{j}$ represent the $j$-th input sample and its corresponding label, respectively.
@@ 4 Method
This section introduces GUIDE - a novel method designed to mitigate catastrophic forgetting in a classifier trained in a generative replay scenario with a diffusion model. We benefit from the classifier trained on the current task, referred to as the current classifier, to guide the diffusion model during the generation of rehearsal examples from preceding tasks. With our method, we can generate rehearsal examples close to the classifier’s decision boundary, making them highly valuable to counteract the classifier’s forgetting in class-incremental learning [44, 26].
@@ 4.1 Intuition and rationale behind GUIDE
Before moving to the continual-learning setup, we demonstrate the effect of guiding the diffusion model towards classes not included in its training dataset. To that end, we propose a simplified scenario in which we employ the unconditional diffusion model $\epsilon_{\theta}(\mathbf{x}_{t},t)$ trained exclusively on the goldfish and tiger shark classes from the ImageNet100 dataset, along with a classifier $f_{\phi}(y|\mathbf{x})$ trained on entire ImageNet dataset.
Given the diffusion model’s unconditional nature, we use the classifier to steer the denoising process toward either the goldfish or tiger shark class, represented as $c_{1}$. Simultaneously, we add another guiding signal from the same classifier $f_{\phi}(y|\mathbf{x})$ towards one of the classes from the ImageNet dataset different than goldfish and tiger shark, denoted as $c_{2}$. Formally, we modify the unconditional diffusion sampling process as follows:
${\hat{\epsilon}}_{\theta}(\mathbf{x}_{t},t)=\epsilon_{\theta}(\mathbf{x}_{t},t)+s_{1}\nabla_{\mathbf{x}_{t}}{\ell}\left(f_{\phi}(y|\hat{\mathbf{z}}_{0}(\mathbf{x}_{t})),c_{1}\right)+s_{2}\nabla_{\mathbf{x}_{t}}{\ell}\left(f_{\phi}(y|\hat{\mathbf{z}}_{0}(\mathbf{x}_{t})),c_{2}\right).$ (8)
In Fig. 2, we present generated samples, with goldfish generations in the upper row and tiger sharks in the bottom one. Additionally, we present more samples generated using this technique in Appendix H.
By integrating guidance to both $c_{1}$ and $c_{2}$ classes, we generate samples from the diffusion model’s training data distribution – all of the examples are either goldfishes or sharks but with visible features from classes $c_{2}$, which are unknown to the diffusion model (e.g., oblong shape of baguette, hay or snow in the background).
In this work, we propose to use this observation in the continual training of a classifier, with the distinction that in GUIDE, we only utilize guidance toward unknown classes since we sample from a class-conditional diffusion model. Thus, we eliminate the need the need for guidance toward classes from the diffusion model’s training set.
Figure 2: Samples from the unconditional diffusion model trained only on goldfish and tiger shark classes from the ImageNet100 dataset. In the upper row, we present the samples guided to the goldfish class, while in the bottom row, to the tiger shark class. At the same time, the classifier guides the denoising process toward the class depicted above each figure that was not included in the training set of the diffusion model. For reference, in the leftmost column, we present samples generated without guidance toward any unknown class, setting $s_{2}=0$. In every other column, we set both $s_{1}$ and $s_{2}$ to $10$. We obtain samples from the desired class with observable features of classes unknown to the diffusion model, such as the color, background, or shape.